OpenStudy (anonymous):

Help with two questions, medal!! 1) A megabyte is 2^20. A terabyte is 2^40. Using the exponent rule for division, find how many megabytes are in a terabyte. 2) Explain how to convert 2.7 × 10^4 from scientific notation to standard form, then give the answer in standard form.

OpenStudy (usukidoll):

I do know that 10 to a positive exponent will have results like 10 100 1000 10^1 = 10 10^2 = 100 10^3 = 1000 10^4 = 10000 2.7 x 10,000 I think you have to move the decimal to the right by 4 places. for the first one ... 2^40/2 ^ 20 ====> 2^(40-20) = 2^20... I think :/ but it did say exponent rule for division.

OpenStudy (anonymous):

I just seen that you replied, but I got the first question I'm just having some trouble with the second one. Thanks for the help though!